Tohban Report 2013-02-20
From RHESSI Wiki
|Start Date:||13 Feb 2013|
|End Date:||20 Feb 2013|
|Next Tohban:||Hazel Bain|
|List all reports|
Some activity, with an M-class flare.
How many GOES flares occurred?
Flares above B, C, M, X class were 17 6 1 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 11 5 1 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 8 4 1 0
There were RHESSI flares/GOES flares 134 / 24 over the time range 12-Feb-13 19-Feb-13
No memory issues. The SSR is emptying at the end of the daily pass set, with a maximum of about 25%.
The spacecraft parameters are nominal. No HV commanding this week but a few commanding to change D2 threshold values (see below). We may need a spin-up next week.
No gaps in monitor rates.
We lowered the front slow threshold on detector 2, but there was no noticeable change in front segment counts after the changes were made. The detector remains unsegmented.
DOY 044, 22:38 UT, lowered from 0x88 to 0x58, no change in count rates DOY 045, 19:08 UT, lowered from 0x58 to 0x48, gradual increase with fluctuations DOY 045, 20:47 UT, raised from 0x48 to 0x50, a little decrease
Looking at the count rate plots, the fluctuations we observed at the time of commanding seem to have been natural. Energy spectra generated from the detector 2 now show a bump in the 15-25 keV range, but it is not certain whether this is due to the threshold changes. There could have been some effect from light-leak. From next week, we will try lowering the threshold until we really see significant noise.
A new page "DIB tables" is added to the Wiki site under the page "Tohban detector updates".
At the Ops meeting, Säm Krucker pointed out that the SAS calibration may be off. Martin Fivian will be looking at it.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||+/- 4 minutes|
|Night time data (rears)||Taking data at nights|
|Require extra passes?||No|
|Requirement for moving pointer?||No|